OpenStudy (anonymous):

Find the sum of the infinite series. 1.2 + 0.12 + 0.012 +

OpenStudy (jhannybean):

this is a geometric series where the ratio is 0.1. The formula for a geoemtric series is \[\large \sum_{n=1}^{\infty}a_{1}r^{n-1}\]\[\large a_{1}= 1.2 \ ,\ r= 0.1\] \[\large |r|<1 \ \text{ series converges}\]

OpenStudy (jhannybean):

And if the series converges, then you find the sum \[\large s_{n}= \frac{a}{1-r}\]

OpenStudy (jhannybean):

So first, does your series converge? If so, plug your values into the formula for the sum, and there you go.

OpenStudy (anonymous):

how do you know if it converges?

OpenStudy (jhannybean):

if |r|<1

OpenStudy (jhannybean):

what is your ratio in this series?

OpenStudy (anonymous):

0.1 right?

OpenStudy (jhannybean):

Yes, is |0.1| < 1?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

ok, so now we know the series converges. To find the sum, plug a = 1.2 and r = 0.1 into s(n) = (a)/ (1-r)

OpenStudy (anonymous):

1.1?

OpenStudy (jhannybean):

\[\large \frac{1.2}{1-(0.1)}=\frac{1.2}{0.9}=?\]

OpenStudy (anonymous):

1.3333

OpenStudy (jhannybean):

good job